# Cybersecurity AI: Humanoid Robots as Attack Vectors

### v2

@@ Abstract
We present a systematic security assessment of the Unitree G1 humanoid showing it operates simultaneously as a covert surveillance node and can be purposed as an active cyber operations platform. Partial reverse engineering of Unitree’s proprietary FMX encryption reveal a static Blowfish-ECB layer and a predictable LCG mask—enabled inspection of the system’s otherwise sophisticated security architecture, the most mature we have observed in commercial robotics. Two empirical case studies expose the critical risk of this humanoid robot: (a) the robot functions as a trojan horse, continuously exfiltrating multi-modal sensor and service-state telemetry to 43.175.228.18:17883 and 43.175.229.18:17883 every 300 seconds without operator notice, creating violations of GDPR Articles 6 and 13; (b) a resident Cybersecurity AI (CAI) agent can pivot from reconnaissance to offensive preparation against any target, such as the manufacturer’s cloud control plane, demonstrating escalation from passive monitoring to active counter-operations. These findings argue for adaptive CAI-powered defenses as humanoids move into critical infrastructure, contributing the empirical evidence needed to shape future security standards for physical–cyber convergence systems.
@@ 1 Introduction
This technical brief summarizes the findings from Alias Robotics’ security assessment of the Unitree G1 humanoid robot conducted in September 2025. Refer to the full report for more details. [^1] Full technical report: Mayoral-Vilches, V., ”The Cybersecurity of a Humanoid Robot: An Early Study on the Cybersecurity of Humanoid Robots via the Unitree G1”, Alias Robotics, September 2025. Our methodology builds upon established prior robotics security assessment practices [1, 2, 3, 4, 5, 6, 7, 8, 9] combining static firmware analysis, binary reverse engineering of the 9.2 MB master_service orchestrator, runtime network traffic analysis to systematically evaluate the platform’s security architecture and more contemporary research on Cybersecurity AI [10, 11, 12, 13]. Four critical findings emerged:
@@
The remainder of the brief walks through the platform’s anatomy, the FMX cryptanalytic analysis, the telemetry pipeline, and two validated attack vectors before closing with what the dual-threat reality means.
@@ 2.1  Hardware Baseline
Hardware analysis reveals the G1’s compute platform centers on a Rockchip RK3588 SoC (quad-core Cortex-A76 @ 2.4GHz + quad-core Cortex-A55 @ 1.8GHz) with 8GB LPDDR4X RAM and 32GB eMMC storage. Physical security weaknesses include exposed JST debug connectors, unpopulated JTAG pads, and accessible UART interfaces operating at 115200 baud. The sensor array comprises Intel RealSense D435i depth cameras (1280×720 @ 30fps depth, 1920×1080 @ 30fps RGB), dual MEMS microphones, 9-axis IMU, and GNSS receiver—all publishing to DDS topics on the local network. Some of these sensors are susceptible to established attack vectors including optical channel exploitation [15, 16] and acoustic manipulation [17].
@@ 2.3  Communication Surfaces
  • BLE/Wi-Fi links via upper_bluetooth and chat_go for mobile or app-based control.
@@ 5.2  Case Study 2: Weaponized Cybersecurity AI Platform
  • Recon. Enumerated live connections and identified MQTT/WebSocket/WebRTC endpoints reachable from inside the robot.
  • Exploitation rehearsal. Attempted broker logins using the recovered credentials (anonymous access rejected, certificate-authenticated sessions prepped but withheld during ethical testing), demonstrating the feasibility of command injection and telemetry spoofing.
  • Attack planning. Produced an attack surface matrix covering MQTT topic abuse, WebRTC stream hijack, and OTA manipulation pathways.

### v3

@@ Abstract
We present a systematic security assessment of the Unitree G1 humanoid showing it operates simultaneously as a covert surveillance node and can be purposed as an active cyber operations platform. Initial access can be achieved by exploiting the BLE provisioning protocol which contains a critical command injection vulnerability allowing root access via malformed Wi-Fi credentials, exploitable using hardcoded AES keys shared across all units. Partial reverse engineering of Unitree’s proprietary FMX encryption reveal a static Blowfish-ECB layer and a predictable LCG mask—enabled inspection of the system’s otherwise sophisticated security architecture, the most mature we have observed in commercial robotics. Two empirical case studies expose the critical risk of this humanoid robot: (a) the robot functions as a trojan horse, continuously exfiltrating multi-modal sensor and service-state telemetry to 43.175.228.18:17883 and 43.175.229.18:17883 every 300 seconds without operator notice, creating violations of GDPR Articles 6 and 13; (b) a resident Cybersecurity AI (CAI) agent can pivot from reconnaissance to offensive preparation against any target, such as the manufacturer’s cloud control plane, demonstrating escalation from passive monitoring to active counter-operations. These findings argue for adaptive CAI-powered defenses as humanoids move into critical infrastructure, contributing the empirical evidence needed to shape future security standards for physical–cyber convergence systems.
@@ 1 Introduction
This technical brief summarizes the findings from Alias Robotics’ security assessment in combination with researchers Andreas Makris3 and Kevin Finisterre2 independent findings [^1] Refer to https://github.com/Bin4ry/UniPwn for Andreas Makris’ and Kevin Finisterre’s exploitation of this vulnerability. of the Unitree G1 humanoid robot. Refer to the full report [1] and for more details. Our methodology builds upon established prior robotics security assessment practices [2, 3, 4, 5, 6, 7, 8, 9, 10] combining static firmware analysis, binary reverse engineering of the 9.2 MB master_service orchestrator, runtime network traffic analysis, and BLE protocol exploitation to systematically evaluate the platform’s security architecture and more contemporary research on Cybersecurity AI [11, 12, 13, 14]. Five critical findings emerged:
@@
The remainder of the brief walks through the platform’s anatomy, BLE provisioning vulnerabilities, the FMX cryptanalytic analysis, the telemetry pipeline, and validated attack vectors before closing with what the dual-threat reality means.
@@ 2.1  Hardware Baseline
Hardware analysis reveals the G1’s compute platform centers on a Rockchip RK3588 SoC (quad-core Cortex-A76 @ 2.4GHz + quad-core Cortex-A55 @ 1.8GHz) with 8GB LPDDR4X RAM and 32GB eMMC storage. Physical security weaknesses include exposed JST debug connectors, unpopulated JTAG pads, and accessible UART interfaces operating at 115200 baud. The sensor array comprises Intel RealSense D435i depth cameras (1280×720 @ 30fps depth, 1920×1080 @ 30fps RGB), dual MEMS microphones, 9-axis IMU, and GNSS receiver—all publishing to DDS topics on the local network. Some of these sensors are susceptible to established attack vectors including optical channel exploitation [16] and acoustic manipulation [17].
2.2  BLE Provisioning Vulnerability
[…]
@@ 2.4  Communication Surfaces
  • BLE/Wi-Fi links via upper_bluetooth and chat_go for mobile or app-based control. The BLE channel specifically exposes the command injection vulnerability.
@@ 5.2  Case Study 2: Weaponized Cybersecurity AI Platform
  • Recon. Enumerated live connections and identified MQTT/WebSocket/WebRTC endpoints reachable from inside the robot. Additionally confirmed the BLE provisioning vulnerability allows trivial initial compromise using hardcoded AES keys.
  • Exploitation rehearsal. Attempted broker logins using the recovered credentials (anonymous access rejected, certificate-authenticated sessions prepped but withheld during ethical testing), demonstrating the feasibility of command injection via both BLE provisioning and direct telemetry spoofing.
  • Attack planning. Produced an attack surface matrix covering BLE-based command injection, MQTT topic abuse, WebRTC stream hijack, and OTA manipulation pathways.